calculate;the unemployment rate,labour force participation rate,absorption rate,given the following statistics south africa reported the following data for 2012.labour force.17948000.employment,13422000,working age population 32786000
1
Expert's answer
2019-07-29T08:03:56-0400

Let unemployment rate=A

Labour force participation rate=B

Labour force absorption rate=C

Labour force=a=17948000

employment=b=13422000

Working age population=c=32786000

Then,

"A=\frac{(a-b)}{a}\times 100"

"A=\frac{(17948000-13422000)}{17948000}\times 100"

A=25.21%

"B=\frac{a}{c}\times 100", "B=\frac{17948000}{32786000}\times 100"

B= 54.7%


"C=\frac{b}{c}\times 100"

"C=\frac{13422000}{32786000}\times 100"

C=40.9%